# Critical temperatures of two dimensional magnets beyond linear spin wave theory: application to CrI3, MPS3 (M=Ni, Mn, Fe) and CrSBr

### v1

@@ I Introduction
Since the discovery of ferromagnetic order in monolayer CrI3 [1] there has been a increasing interest in magnetic two-dimensional materials. In addition to the fundamental physical aspects associated with two-dimensional (2D) magnetism, a wide variety of possible spintronics applications have been proposed [2, 3]. Compared to three-dimensional (3D) magnetic materials, a major difference is the crucial role played by magnetic anisotropy. The Mermin-Wagner (MW) theorem states that continuous symmetries cannot be spontaneously broken at finite temperature in 2D systems with short-range interactions [4] and this implies that order requires explicit symmetry breaking by magnetic anisotropy. While one may still obtain a reliable description of the magnetic interactions in terms of Heisenberg models, well-established methodology for calculating magnetic properties of such models may break down in 2D. For example, conventional Weiss mean field theory predicts a finite critical temperature in any isotropic model - irrespective of the dimensionality of the lattice - and thus fails rather dramatically in 2D. On the other hand, correlation effects can be accounted for by applying classical Monte Carlo (MC) simulations, which are commonly used to estimate critical temperatures [5, 6, 7, 8]. However, such calculations are often computationally demanding and accuracy is typically limited by the classical treatment [9]. Finally, one may try to obtain the thermal properties directly from the quantum mechanical excitations, which can be calculated by various approximate methods. In the case of long range order, the fundamental magnetic excitations of the Heisenberg model will be comprised by magnons, and while single-magnon energies are easily calculated, magnon-magnon interactions have to be approximated at the mean-field level. This leads to a temperature-dependent renormalization of magnon energies, which can be calculated in either the Holstein Primakoff (HP) bosonization scheme or in the Green’s function-based Random Phase Approximation (RPA) [10, 11, 12]. These methods have recently been developed and benchmarked against experimental critical temperatures in 3D materials [9], but due to the importance of magnetic anisotropy in 2D, the application to 2D magnetism calls for a study of its own.
@@
At the level of two-spin interactions, spin-orbit coupling may give rise to either anisotropic exchange (two-site) interactions or single-ion (one-site) anisotropy. The former case can be handled rather generally by introducing an exchange tensor describing the magnetic interaction between two sites and is straightforward to include in both HP [13] and RPA [9]. The single-ion anisotropy may be included in these approaches as well, but the associated mean-field approximation becomes much more subtle because the two spin operators do not commute at the same site. For example, the generalized single-$Q$ method of Ref. [9], can handle uniaxial, biaxial and triaxial single-ion anisotropies in the Heisenberg model, but RPA fails to predict the irrelevance of single-ion anisotropy for the case of $S=1/2$ systems. The HP method (with a mean-field treatment of two-magnon interactions) correctly captures The vanishing spin-wave gap (irrelevance of single-ion anisotropy) for uniaxial ferromagnets at $T=0$, but fails to do so at elevated temperatures as well as tri-axial systems at $T=0$. An alternative to RPA was proposed by Herbert Callen in 1963 [14] and will be referred to as Callen Decoupling (CD) in the following. Callen’s scheme comprises an alternative mean-field decoupling of the equation of motion, which may be applied to the calculations of critical temperatures. We have recently shown that in this respect the method performs better than RPA in the classical limit (relevant for itinerant systems), but not in the quantum regime [15]. However, CD gives the correct behavior for the single-ion anisotropy contribution to the magnon dispersion (irrelevance of single-ion anisotropy and vanishing spin-wave gap for $S=1/2$ systems), and it has been shown that augmenting the RPA method for exchange interactions with CD for the single-ion anisotropy (RPA+CD) gives excellent agreement with exact results for ferromagnets [15, 16]. The CD has been developed for simple anti-ferromagnets [17, 18] as well, but a general approach for arbitrary single-$Q$ states seems to be lacking in the literature.
@@ II Theory
where $\mathbf{S}_{ai}$ is the spin operator for site $a$ in unit cell $i$ and $\mathrm{J}_{abij}$ is the exchange tensor that couples spins at $i,a$ and $j,b$. Written in this form the model includes single-ion anisotropy represented by the terms with $a=b$ and $i=j$. In the following it will be practical to consider such terms separately and we write them as
@@
with $\mathrm{K}_{ai}=\mathrm{J}_{aaii}/2$ being the single-ion anisotropy tensor. It should be noted that the magnon dispersion must be invariant under the addition of an isotropic term to the single-ion anisotropy tensor: $\mathrm{K}_{ai}^{\mu\nu}\rightarrow\mathrm{K}_{ai}^{\mu\nu}+\alpha\delta^{\mu\nu}$. This is due to the fact that $\mathbf{S}_{ai}\cdot\mathbf{S}_{ai}$ is a constant and will be referred to as isotropic invariance. This freedom will be utilized to enforce certain components of the tensor to vanish in the following. It is also clear that the single-ion anisotropy term becomes a constant (and thus irrelevant) for any $S=1/2$ system.
@@
where $\hat{\mathbf{n}}$ is the global rotation axis for the single-$Q$ spiral state, $\mathbf{S}^{\prime}_{ai}$ are spins in the coordinates that are independent of unit cells and $\mathbf{S}^{\prime\prime}_{ai}$ are spins in the locally ferromagnetic frame independent of sites [13, 9]. Introducing two new vectors $\mathbf{u}_{a}$ and $\mathbf{v}_{a}$
@@ II.1 Holstein-Primakoff bosonization
In the HP approach we express the spin operators in terms of bosonic operators, which ensure that the commutators between spin operators are conserved. In the local reference frame, the ground state spin direction is always along $z$ and we thus apply the local Holstein-Primakoff transformation [10, 13],
@@
$\quad+\mathbf{u}^{*T}_{a}\mathrm{K}_{a}\mathbf{u}_{a}(2a_{ai}a_{ai}a^{\dagger}_{ai}a^{\dagger}_{ai})$
@@
$\quad+\mathbf{u}_{a}^{T}\mathrm{K}_{a}\mathbf{u}_{a}(2a^{\dagger}_{ai}a^{\dagger}_{ai}a^{\dagger}_{ai}a_{ai}+a^{\dagger}_{ai}a^{\dagger}_{ai}\Bigr),$ (17)
where $N$ is the number of unit cells . We note that normal ordering of the quartic terms generates more quadratic terms that were not taken into account in $\mathcal{H}_{\mathrm{K}}^{2}$ and will affect the zero-temperature dispersion. The bosonic operators are then Fourier transformed:
@@
$\mathrm{A}^{\mathrm{HP}}_{a}$ $=S_{a}\mathbf{u}_{a}^{T}\mathrm{K}_{a}\mathbf{u}^{*}_{a}\Big(1-\frac{1}{S_{a}}(1+2\Phi_{a})\Big),$ (22)
@@
and the magnon occupation $\Phi_{a}$ is defined in Eq. (90). The terms in brackets $\frac{1}{S_{a}}(\dots)$ arise from the quartic Hamiltonian (17) and breaks the isotropic invariance of the $\mathrm{K}_{a}$ tensor. The magnon energies are obtained by (para)diagonalizing Eq. (20) [13, 9]. The sublattice magnetization is then calculated as
@@
$\Phi^{0}$ is the saturation magnon number, obtained by evaluating Eq. (90) at $T=0$. Note that all the site indices have been dropped as we are considering a simple Bravais lattice. The magnon gap is defined as
$\Delta=\omega_{\mathbf{q}=0}$ (29)
@@
where $K=\mathrm{K}^{zz}-\mathrm{K}^{xx}$. We see that the magnon gap vanishes in the case of $S=1/2$, which is the expected result. However, at finite temperatures this no longer holds and for tri-axial systems the magnon dispersion will acquire a dependence on the single-ion anisotropy at $T=0$ for $S=1/2$ and may even lead to instabilities (negative magnon energies). The basic reason for this failure is the truncation of $S^{{\prime\prime}\pm}_{ai}$ in Eqs. (10)-(11), which becomes inaccurate if the anisotropy tensor couples spins that are not along the local $z$-axis.
@@ II.2 Green’s function method
The RPA Green’s function approach has previously been shown to yield rather accurate predictions for critical temperatures in single-$Q$ bulk systems [9]. Here, we extend the RPA to include single-ion anisotropy. The method is then shown to fail rather strikingly in the case of $S=1/2$ and is expected to overestimate spinwave gaps in general. We then proceed with the Callen decoupling applied to single-ion anisotropies and show that the failure of RPA is largely remedied in this approach.
@@ II.2.1 Random Phase Approximation
Here $\mathrm{H}^{\mathrm{RPA}}_{\mathbf{q}}$ contains the contribution from the exchange tensor, which has been derived previously [9] and is stated in appendix B. The single-ion anisotropy contribution, $\mathrm{H}^{\mathrm{RPA}}_{\mathrm{K}}$ is given by
@@
Unlike HP, the isotropic invariance of the tensor $\mathrm{K}_{a}$ is preserved in RPA. This is easily seen by noting that $\mathbf{u}^{T}_{a}\mathbf{u}^{*}_{a}=2$, $\mathbf{u}^{T}_{a}\mathbf{u}_{a}=0$, and $\mathbf{v}^{T}_{a}\mathbf{v}_{a}=1$.
@@
$Y^{\mathrm{RPA}}_{\mathbf{q}}$ $=\frac{\mathbf{u}^{T}(\mathrm{J}^{\prime}_{\mathbf{q}}+2\mathrm{K})\mathbf{u}}{2}.$ (45)
It can be observed that RPA predicts a finite gap for $S=1/2$ and fails to capture the irrelevance of singe-ion anisotropy in this case. This is a serious flaw of the method, which we address by turning to the CD scheme for the single-ion anisotropy.
@@ II.2.2 Callen decoupling
As RPA assumes the fluctuations in $S^{{\prime\prime}z}_{ai}$ to be small, the large fluctuations that would become relevant at higher temperatures are disregarded a priori. The CD attempts to generalize the RPA mean-field approach in such a way that fluctuations in $S^{{\prime\prime}z}_{ai}$ become important at high temperatures [14, 17]. However, comparing results with quantum Monte Carlo simulations for $S=1/2$ ferromagnets reveals that the method actually yields worse results than RPA [16]. For higher spins it is expected to become more accurate and for itinerant systems (treated in the classical limit) the CD was found to be more accurate than RPA [15]. Moreover, it has been shown that the CD predicts much better results than RPA when it is solely applied to single-ion anisotropy in ferromagnets [15, 16]. One reason for this is that it predicts the correct lack of order in $S=1/2$ and is not subject to the overestimation of spinwave gaps inherent to RPA. As the failure of RPA for single-ion anisotropy is not limited to ferromagnets, we derive a single-$Q$ generalization of the CD for single-ion anisotropy. Combined with the RPA Hamiltonian for the exchange thus yields the hybrid RPA+CD approach.
@@
and the parameter $\alpha$ can be chosen as $\frac{\langle S_{a}^{{\prime\prime}z}\rangle}{2S_{a}^{2}}$ which will interpolate between 1 and 0 for $T=0$ and $T=T_{\mathrm{C}}$ respectively. Applying this to Eq. (50), the CD single-ion anisotropy Hamiltonian is obtained as
@@
where $\alpha,\beta\in[1,2N_{a}]$, $\mathrm{H}^{\mathrm{RPA}}_{\mathrm{K}}$ is defined in Eq. (39) and $\Phi_{a}$ is the magnon occupation number defined in Eq. (54). The magnon energies within the RPA+CD method can be obtained by solving the eigenvalue equation (36) with the single-ion anisotropy contribution $\mathrm{H^{RPA}_{K}}$ replaced by $\mathrm{H^{CD}_{K}}$.
@@
where $\Phi^{0}$ is defined in Eq. (59). It is observed that for $S=1/2$, this contribution is close to being independent of $\mathrm{K}_{a}$, as the quantity in bracket is close to (exactly for ferromagnets) zero. Importantly, it is RPA+CD that yields the closest agreement to the true behavior (irrelevance) of single-ion anisotropy in $S=1/2$ systems when comparing the three methods considered here. It is also easy to see that the isotropic invariance of RPA is preserved by CD.

### v2

@@ I Introduction
Since the discovery of ferromagnetic order in monolayer CrI3 [Huang2017] there has been a increasing interest in magnetic two-dimensional materials. In addition to the fundamental physical aspects associated with two-dimensional (2D) magnetism, a wide variety of possible spintronics applications have been proposed [Burch2018, Wang2022-rev]. Compared to three-dimensional (3D) magnetic materials, a major difference is the crucial role played by magnetic anisotropy. The Mermin-Wagner (MW) theorem states that continuous symmetries cannot be spontaneously broken at finite temperature in 2D systems with short-range interactions [MW1966] and this implies that order requires explicit symmetry breaking by magnetic anisotropy. While one may still obtain a reliable description of the magnetic interactions in terms of Heisenberg models, well-established methodology for calculating magnetic properties of such models may break down in 2D. For example, conventional Weiss mean field theory predicts a finite critical temperature in any isotropic model - irrespective of the dimensionality of the lattice - and thus fails rather dramatically in 2D. On the other hand, correlation effects can be accounted for by applying classical Monte Carlo (MC) simulations, which are commonly used to estimate critical temperatures [Torelli2019, Strandburg-mc1992, Xiaobo-mc2019, Xue-mc2022]. However, such calculations are often computationally demanding and accuracy is typically limited by the classical treatment [Helimag3d2025]. Finally, one may try to obtain the thermal properties directly from the quantum mechanical excitations, which can be calculated by various approximate methods. In the case of long range order, the fundamental magnetic excitations of the Heisenberg model will be comprised by magnons, and while single-magnon energies are easily calculated, magnon-magnon interactions have to be approximated at the mean-field level. This leads to a temperature-dependent renormalization of magnon energies, which can be calculated in either the Holstein Primakoff (HP) bosonization scheme or in the Green’s function-based Random Phase Approximation (RPA) [HP1940, tyablikov1959, Tahir1962]. These methods have recently been developed and benchmarked against experimental critical temperatures in 3D materials [Helimag3d2025], but due to the importance of magnetic anisotropy in 2D, the application to 2D magnetism calls for a study of its own.
@@
At the level of two-spin interactions, spin-orbit coupling may give rise to either anisotropic exchange (two-site) interactions or single-ion (one-site) anisotropy. The former case can be handled rather generally by introducing an exchange tensor describing the magnetic interaction between two sites and is straightforward to include in both HP [Toth2015] and RPA [Helimag3d2025]. The single-ion anisotropy may be included in these approaches as well, but the associated mean-field approximation becomes much more subtle because the two spin operators do not commute at the same site. For example, the generalized single-$Q$ method of Ref. [Helimag3d2025], can handle uniaxial, biaxial and triaxial single-ion anisotropies in the Heisenberg model, but RPA fails to predict the irrelevance of single-ion anisotropy for the case of $S=1/2$ systems. The HP method (with a mean-field treatment of two-magnon interactions) correctly captures the vanishing spin-wave gap (irrelevance of single-ion anisotropy) for uniaxial $S=1/2$ ferromagnets at $T=0$, but fails to do so at elevated temperatures as well as tri-axial systems at $T=0$. An alternative to RPA was proposed by Herbert Callen in 1963 [Callen1963] and will be referred to as Callen Decoupling (CD) in the following. Callen’s scheme comprises an alternative mean-field decoupling of the equation of motion, which may be applied to the calculations of critical temperatures. We have recently shown that in this respect the method performs better than RPA in the classical limit (relevant for itinerant systems), but not in the quantum regime [BeyondRPA2025]. However, CD gives the correct behavior for the single-ion anisotropy contribution to the magnon dispersion (irrelevance of single-ion anisotropy and vanishing spin-wave gap for $S=1/2$ systems), and it has been shown that augmenting the RPA method for exchange interactions with CD for the single-ion anisotropy (RPA+CD) gives excellent agreement with exact results for ferromagnets [BeyondRPA2025, Frobich2006]. The CD has been developed for simple anti-ferromagnets [Callen1964-afm, Liu-CD1967] as well, but a general approach for arbitrary single-$Q$ states seems to be lacking in the literature.
@@ II Theory
where $\mathbf{S}_{ai}$ is the spin operator for site $a$ in unit cell $i$ and $\mathrm{J}_{abij}$ is the exchange tensor that couples spins at $i,a$ and $j,b$. We will denote the number of unit cells by $N$ and the number of magnetic sites per unit cell by $N_{a}$. Written in this form the model includes single-ion anisotropy represented by the terms with $a=b$ and $i=j$. In the following it will be practical to consider such terms separately and we write them as
@@
with $\mathrm{K}_{ai}=\mathrm{J}_{aaii}/2$ being the single-ion anisotropy tensor. It should be noted that the magnon dispersion must be invariant under the addition of an isotropic term to the single-ion anisotropy tensor: $\mathrm{K}_{ai}^{\mu\nu}\rightarrow\mathrm{K}_{ai}^{\mu\nu}+\alpha\delta^{\mu\nu}$. This is due to the fact that $\mathbf{S}_{ai}\cdot\mathbf{S}_{ai}$ is a constant and will be referred to as isotropic invariance. The freedom will be utilized to enforce certain components of the tensor to vanish in the following. It is also clear that the single-ion anisotropy term becomes a constant (and thus irrelevant) for any $S=1/2$ system.
@@
where $\hat{\mathbf{n}}$ is the global rotation axis for the single-$Q$ spiral state, $\mathbf{S}^{\prime}_{ai}$ are spins in the coordinates that are independent of unit cells and $\mathbf{S}^{\prime\prime}_{ai}$ are spins in the locally ferromagnetic frame independent of sites [Toth2015, Helimag3d2025]. Introducing two new vectors $\mathbf{u}_{a}$ and $\mathbf{v}_{a}$
@@ II.1 Holstein-Primakoff bosonization
In the HP approach we express the spin operators in terms of bosonic operators, which ensure that the commutators between spin operators are conserved. In the local reference frame, the ground state spin direction is always along $z$ and we thus apply the local Holstein-Primakoff transformation [HP1940, Toth2015],
@@
$\quad+\mathbf{u}^{*T}_{a}\mathrm{K}_{a}\mathbf{u}_{a}(2a^{\dagger}_{ai}a^{\dagger}_{ai}a_{ai}a_{ai}+4a^{\dagger}_{ai}a_{ai})$
@@
$\quad+\mathbf{u}_{a}^{T}\mathrm{K}_{a}\mathbf{u}_{a}(2a^{\dagger}_{ai}a^{\dagger}_{ai}a^{\dagger}_{ai}a_{ai}+a^{\dagger}_{ai}a^{\dagger}_{ai}\Bigr).$ (17)
We note that normal ordering of the quartic terms generates more quadratic terms that were not taken into account in $\mathcal{H}_{\mathrm{K}}^{2}$ and will affect the zero-temperature dispersion. The bosonic operators are then Fourier transformed:
@@
$\mathrm{A}^{\mathrm{HP}}_{a}$ $=S_{a}\mathbf{u}_{a}^{T}\mathrm{K}_{a}\mathbf{u}^{*}_{a}\Big(1-\frac{1}{2S_{a}}(1+4\Phi_{a})\Big),$ (22)
@@
and the magnon occupation $\Phi_{a}$ is defined in Eq. (91). The terms in brackets $\frac{1}{S_{a}}(\dots)$ arise from the quartic Hamiltonian (17). The isotropic invariance is preserved by these terms, which is easily seen by noting that $\mathbf{u}^{T}_{a}\mathbf{u}^{*}_{a}=2$, $\mathbf{u}^{T}_{a}\mathbf{u}_{a}=0$, and $\mathbf{v}^{T}_{a}\mathbf{v}_{a}=1$. The magnon energies are obtained by (para)diagonalizing Eq. (20) [Toth2015, Helimag3d2025]. The sublattice magnetization is then calculated as
@@
$\Phi^{0}$ is the saturation magnon number, obtained by evaluating Eq. (91) at $T=0$ and $\mathrm{J}^{\prime}_{\mathbf{q}}$ is defined in Eq. (85). Note that all the site indices have been dropped as we are considering a simple Bravais lattice. The magnon gap is defined as
$\Delta=\omega^{\textrm{HP}}_{\mathbf{q}=0}$ (29)
@@
where $K=\mathrm{K}^{zz}-\mathrm{K}^{xx}$. This is obtained by evaluating Eq. (29) with $\mathbf{u}=[1,i,0]$ and $\mathbf{v}=[0,0,1]$ (ground state magnetization along the $z$-axis). We see that the magnon gap vanishes in the case of $S=1/2$, which is the expected result. However, at finite temperatures this no longer holds and for tri-axial systems the magnon dispersion will acquire a dependence on the single-ion anisotropy at $T=0$ for $S=1/2$ and may even lead to instabilities (negative magnon energies). The basic reason for this failure is the truncation of $S^{{\prime\prime}\pm}_{ai}$ in Eqs. (10)-(11), which becomes inaccurate if the anisotropy tensor couples spins that are not along the local $z$-axis.
@@ II.2 Green’s function method
The RPA Green’s function approach has previously been shown to yield rather accurate predictions for critical temperatures in single-$Q$ bulk systems [Helimag3d2025]. Here, we extend the RPA to include single-ion anisotropy. The method is then shown to fail rather strikingly in the case of $S=1/2$ and is expected to overestimate spinwave gaps in general. We then proceed with the Callen decoupling applied to single-ion anisotropies and show that the failure of RPA is largely remedied in this approach.
@@ II.2.1 Random Phase Approximation
Here $\mathrm{H}^{\mathrm{RPA}}_{\mathbf{q}}$ contains the contribution from the exchange tensor, which has been derived previously [Helimag3d2025] and is stated in appendix B. The single-ion anisotropy contribution, $\mathrm{H}^{\mathrm{RPA}}_{\mathrm{K}}$ is given by
@@
It is straightforward to verify that isotropic invariance is conserved by $\mathrm{H}^{\mathrm{RPA}}_{\mathrm{K}}$.
@@
$Y^{\mathrm{RPA}}_{\mathbf{q}}$ $=\frac{\mathbf{u}^{T}(\mathrm{J}^{\prime}_{\mathbf{q}}+2\mathrm{K})\mathbf{u}}{2},$ (45)
where $\mathrm{J}^{\prime}_{\mathbf{q}}$ is defined in Eq. (85). It can be observed that RPA predicts a finite gap for $S=1/2$ and fails to capture the irrelevance of singe-ion anisotropy in this case. This is a serious flaw of the method, which we address by turning to the CD scheme for the single-ion anisotropy.
@@ II.2.2 Callen decoupling
As RPA assumes the fluctuations in $S^{{\prime\prime}z}_{ai}$ to be small, the large fluctuations that would become relevant at higher temperatures are disregarded a priori. The CD attempts to generalize the RPA mean-field approach in such a way that fluctuations in $S^{{\prime\prime}z}_{ai}$ become important at high temperatures [Callen1963, Callen1964-afm]. However, comparing results with quantum Monte Carlo simulations for $S=1/2$ ferromagnets reveals that the method actually yields worse results than RPA [Frobich2006]. For higher spins it is expected to become more accurate and for itinerant systems (treated in the classical limit) the CD was found to be more accurate than RPA [BeyondRPA2025]. Moreover, it has been shown that the CD predicts much better results than RPA when it is solely applied to single-ion anisotropy in ferromagnets [BeyondRPA2025, Frobich2006]. One reason for this is that it predicts the correct lack of order in $S=1/2$ and is not subject to the overestimation of spinwave gaps inherent to RPA. As the failure of RPA for single-ion anisotropy is not limited to ferromagnets, we derive a single-$Q$ generalization of the CD for single-ion anisotropy. Combined with the RPA Hamiltonian for the exchange thus yields the hybrid RPA+CD approach.
@@
and the parameter $\alpha$ can be chosen as $\frac{\langle S_{a}^{{\prime\prime}z}\rangle}{2S_{a}^{2}}$ satisfying the physical requirements discussed in Ref. [Callen1963]. Applying this to Eq. (50), the CD single-ion anisotropy Hamiltonian is obtained as
@@
where $\alpha,\beta\in[1,2N_{a}]$, $\mathrm{H}^{\mathrm{RPA}}_{\mathrm{K}}$ is defined in Eq. (39) and $\Phi_{a}$ is the magnon occupation number defined in Eq. (55). Similar to RPA, the magnon energies within the RPA+CD method can be obtained by solving the eigenvalue equation
$\big((\omega+i\epsilon)\mathrm{I}-\big(\mathrm{H}^{\mathrm{RPA}}_{\mathbf{q}}+\mathrm{H}^{\mathrm{CD}}_{\mathrm{K}}\big)\big)\mathrm{G}_{\mathbf{q}}(\omega)=\zeta,$ (52)
with $\mathrm{H}^{\mathrm{RPA}}_{\mathbf{q}}$ being the RPA exchange contribution in Eq. (93) and $\mathrm{H}^{\mathrm{CD}}_{\mathbf{q}}$ is given by Eq. (51).
@@
where $\Phi^{0}$ is defined in Eq. (60). It is observed that for $S=1/2$, this contribution is close to being independent of $\mathrm{K}_{a}$, as the quantity in bracket is close to (exactly for ferromagnets) zero. Importantly, it is RPA+CD that yields the closest agreement to the true behavior (irrelevance) of single-ion anisotropy in $S=1/2$ systems when comparing the three methods considered here.